OpenStudy (anonymous):

The traffic cone is 19 inches tall & has a radius of 5 inches. Find the lateral area.

OpenStudy (anonymous):

A circular pillar candle is 2.8 inches wide & 6 inches tall. Find the VOLUME of the candle.

OpenStudy (anonymous):

A tea bag is shaped like a regular square pyramid. Each leg of the base is 4 cm, and the height is 5 cm. Find the VOLUME of the tea bag.

OpenStudy (anonymous):

One question at a time, please... The surface area for a cone is \[A=\pi r^2+\pi r\sqrt{r^2+h^2}\] (if I'm remembering correctly, that is...) The total area is the sum of the lateral area and the area of the circular base. The area of the circular base is \(\pi r^2\), so the lateral area must be \(\pi r\sqrt{r^2+h^2}\).

OpenStudy (anonymous):

The candle is a "circular pillar," which I would take to mean it is cylindrical. The volume of a cylinder is \[V=\pi r^2h\] However, you're given the width, which I would interpret as the diameter, which is twice the length of the radius, so in fact you have \(r=\dfrac{2.8}{2}=1.4\) inches.

OpenStudy (anonymous):

The volume of a square-base regular pyramid is \[V=\frac{1}{3}s^2h\] where \(s\) is the length of each side of the base.

OpenStudy (anonymous):

A regulation softball has a circumference of 12 inches. What is the volume of the softball